### v1

# CP Violation of Baryon Decays from $N\pi$ Scatterings

Abstract
There is a long-standing puzzle that the CP violation (CPV) in the baryon systems has never been well established in experiments, while the CPV of mesons have been observed by decades. In this paper, we propose that the CPV of baryon decays can be generated by the scatterings of a nucleon and a pion into some final states, i.e. $N\pi$ scatterings. Benefited by the fruitful data of $N\pi$ scatterings, we can model-independently analyse the strong phases of $b$-baryon decays using the partial wave amplitudes of $N\pi$ scatterings. This makes a great advantage to predict the CPV of baryon decays. We study the processes of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$, $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$. It is found that the total CPV of the above processes in the invariant mass region of $N\pi$ scatterings are at the order of several percent. More importantly, the regional CPV in the Dalitz plot can reach as large as $20\%$. Therefore, we suggest to measure the CPV of these processes, which has a large possibility to achieve the first observation of CPV in the baryon system.
Introduction.— The matter-antimatter asymmetry in the Universe requires the charge-parity (CP) violation in the microscopic particles Sakharov:1967dj . CP violation (CPV) has been observed in the $K$, $B$ and $D$ mesons sequentially during the past 60 years ParticleDataGroup:2022pth . But it has never been well established in any baryon system, which is a long standing puzzle. In the fact that the visible matter of the universe is dominated by baryons, it is very important to study and search for CPV of baryon systems.
In the Standard Model (SM) of particle physics, baryons do not mix with anti-baryons. CPV of baryons happens only in their decays, which is typically direct CPV. It is well known that the CPV always stems from the interference between tree and penguin diagrams in an individual process, such as $\Lambda_{b}^{0}\to p\pi^{-}$ or $pK^{-}$. However, it has not been observed for the CPV of the above processes even at the level of precision of $\mathcal{O}(1\%)$ LHCb:2018fly . In the multi-body decays, there exist a lot of resonances whose interference can also generate CPV. For example, in the three-body decays of $B^{+}\to\pi^{+}\pi^{+}\pi^{-}$ or $B^{+}\to\pi^{+}K^{+}K^{-}$, the interference between $S$-wave and $P$-wave amplitudes, and the $\pi\pi\leftrightarrow K\bar{K}$ rescattering, contribute to large strong phases and thus giant regional CPV of the Dalitz plot, which could be as large as $(50-60)\%$ LHCb:2019jta ; LHCb:2019xmb ; LHCb:2022fpg ; AlvarengaNogueira:2015wpj ; Cheng:2020ipp . Therefore, it deserves to investigate the multi-body decays of $b$-baryons, such as $\Lambda_{b}^{0}\to pK^{-}\pi^{+}\pi^{-}$ and $p\pi^{-}\pi^{+}\pi^{-}$ which have large data samples at LHCb. If there also exists large regional CPV, it is then possible to observe CPV of baryon decays.
There are a lot of $N^{*}$ excited states in the invariant masses of $p\pi$, $p\pi\pi$ and $\Lambda K$. They could provide fruitful effects of interference and possibly large strong phases, thus possibly large regional CPV. However, there exists a problem meanwhile, that the number of excited $N^{*}$ states are too large. For example, there are more than 15 $N^{*}$ states below 2 GeV. They have different quantum numbers of spin and parity. Their masses and widths are usually of large uncertainties. This makes a significantly difficulty for both experimental partial-wave amplitude analysis and theoretical predictions.
Actually, the excited $N^{*}$ states are usually determined in the experiments of $N\pi$ scatterings. Fortunately, scientists have analyzed the data of $N\pi$ scatterings and provided the partial-wave amplitudes for us. This can be easily obtained from the websites, such as the SAID program SAID . This helps us to model-independently calculate the strong phases of $\Lambda_{b}$ decays, regardless of the detailed information of the excited states. It naturally solves the problem of too many resonances and large uncertainties from the excited $N^{*}$ states. The strong phases are usually of the most difficulty and the largest uncertainty in the predictions of CPV in theory. But using the data of scatterings can significantly improve the theoretical precision on the predictions of CPV. For example, the $\pi\pi\leftrightarrow K\bar{K}$ rescattering data leads to a very accurate description of CPV in $B^{\pm}\to K^{\pm}\pi^{+}\pi^{-}$ and $K^{\pm}K^{+}K^{-}$ Garrote:2022uub , and the enhancement of CPV of $D^{0}\to K^{+}K^{-}$ and $\pi^{+}\pi^{-}$ Bediaga:2022sxw . Analogously, the partial-wave amplitudes obtained in the data of $N\pi\to p\pi$ or $p\pi\pi$ scatterings would be helpful to achieve a model-independent determination of strong phases, thus improve the precision of predictions on CPV of $\Lambda_{b}$ decays.
In this Letter, we propose that there might be large regional CPV in the decays of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$, $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$. The mechanism of CPV in these processes is depicted in Fig. 1. In the short-distance weak interaction, $\Lambda_{b}^{0}$ transited into $N\pi$ with an emitted $\pi^{-}$ or $K^{-}$. Then $N\pi$ scatters into $p\pi^{+}\pi^{-}$ or $p\pi^{0}$ via a long-distance strong interaction. It can be model-independently obtained for the partial-wave amplitudes of scatterings of $N\pi\to p\pi^{+}\pi^{-}$ and $p\pi^{0}$ from the data analysis by the SAID program SAID . This is the most important feature of this proposal. The interference effect between the partial-wave amplitudes are different for the tree and penguin operators, which results in the strong phase difference between the tree and penguin amplitudes of $\Lambda_{b}$ decays. The strong phases from $N\pi$ scatterings are usually large. So it is possible to generate large regional CPV.
This is the first proposal for the CPV mechanism of baryon decays stemming from $N\pi$ scatterings. On one hand, the strong phases can be model-independently obtained, so that the theoretical predictions are more reliable. It realizes the remarkable dynamical predictions of CPV in $\Lambda_{b}$ decays with long-distance contributions. On the other hand, we can do the angular analysis since the data of $N\pi$ scatterings provides very comprehensive partial-wave amplitudes. The interference between different partial-wave amplitudes can be revealed in the angular distributions. From a detailed numerical analysis, we predict both the global and regional CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$, in the mass regions of scatterings of $N\pi\to p\pi^{+}\pi^{-}$ and $p\pi^{0}$. It is found that the global CPV of these processes are several percent, while the regional CPV could reach as large as $(20-30)\%$. It could be even larger for the CPV of some moments of $\cos\theta$ in the angular distributions. Some suggestions are then given for the experimental measurements, which are of large possibility for the first observation of baryon CPV.
CPV mechanism from $N\pi$ scatterings.— The CPV mechanism of baryon decays from the $N\pi$ scatterings is depicted in Fig.1.
Figure 1: Illustration of the CPV mechanism driven by $N\pi$ scatterings. The weak phases come from the tree and penguin operators, while the strong phases stem from the $N\pi$ scatterings. The processes are $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ with $h=\pi$ or $K$.
CPV comes from the interference between the figures of Figs. 1 (a) and (b). The weak phases are from tree and penguin operators, while the strong phases stem from the $N\pi$ scatterings except for the effective Wilson coefficients.
The decay amplitudes of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})h^{-}$ and $(p\pi^{0})h^{-}$ undergo two steps, $\mathcal{A}=\mathcal{S}^{1/2}\mathcal{A}_{0}$, given by the decay mechanism of final-state interaction of weak decays of heavy hadrons Suzuki:1999uc ; Cheng:2004ru ; Chua:2007cm . $\mathcal{A}_{0}$ is the short-distance weak decays of $\Lambda_{b}^{0}\to(N\pi)h^{-}$, while $\mathcal{S}^{1/2}$ represents the long-distance scatterings of $N\pi\to p\pi^{+}\pi^{-}$ or $p\pi^{0}$. The CPV mechanism of $N\pi$ scatterings is similar to the $\pi\pi\to K\bar{K}$ rescattering in $B$-meson three-body decays Bediaga:2013ela ; AlvarengaNogueira:2015wpj ; Cheng:2016shb ; Garrote:2022uub . In $B$-meson decays, the rescattering of $\pi\pi\to K\bar{K}$ happens between two channels, like $B^{\pm}\to\pi^{\pm}\pi^{+}\pi^{-}$ and $\pi^{\pm}K^{+}K^{-}$. The $N\pi$ scatterings in $\Lambda_{b}$ decays contribute to one individual process but with different partial waves.
Firstly, $\Lambda_{b}$ decays via a weak interaction at a very short time, with $b\to u\bar{u}q$, $q=d/s$. This is a short-distance dynamics due to the large mass scale of $b$ quark and the weak interaction. In the factorization hypothesis, it can be naively separated into $\Lambda_{b}\to N\pi$ transition form factors and the decay constants of emitted $h^{-}=\pi^{-}/K^{-}$ in both the tree and penguin diagrams. The full expression of the transition form factors of $\Lambda_{b}\to N\pi$ is long and not necessary to be given at the current stage. We will only consider the $\gamma_{\mu}(\gamma_{5})$ terms for convenience. In order to connect with the $N\pi$ scatterings in the second step, the $N\pi$ systems are taken as a whole with different spin-parity spinors. The matrix element of transition from $\Lambda_{b}$ into $N\pi$ is
$\langle N\pi|(\bar{u}b)_{\rm V-A}|\Lambda_{b}\rangle$
$=\bar{u}_{N\pi}^{1/2^{+}}(f_{1}^{1/2^{+}}\gamma_{\mu}-g_{1}^{1/2^{+}}\gamma_{\mu}\gamma_{5}+\cdots)u_{\Lambda_{b}}$
$+\bar{u}_{N\pi}^{1/2^{-}}(-f_{1}^{1/2^{-}}\gamma_{\mu}+g_{1}^{1/2^{-}}\gamma_{\mu}\gamma_{5}+\cdots)u_{\Lambda_{b}}$
$+\bar{u}_{N\pi}^{3/2^{+}}(f_{1}^{3/2^{+}}\gamma_{\mu}-g_{1}^{3/2^{+}}\gamma_{\mu}\gamma_{5}+\cdots)u_{\Lambda_{b}}$
$+\bar{u}_{N\pi}^{3/2^{-}}(-f_{1}^{3/2^{-}}\gamma_{\mu}+g_{1}^{3/2^{-}}\gamma_{\mu}\gamma_{5}+\cdots)u_{\Lambda_{b}}$
++$\cdots$,+ ⋯ , (1)
where the form factors and the spinors are labelled by the spin-parity quantum numbers of the $N\pi$ systems.
The $N\pi$ system produced via short-distance interaction undergoes a long-distance scattering process, i.e. the right-handed parts of Figs. 1 (a) and (b). We consider the scatterings of $N\pi\to p\pi^{+}\pi^{-}$ and $p\pi^{0}$. In practice, $p\pi^{+}\pi^{-}$ is dominated by $\Delta^{++}\pi^{-}$. If necessary, it could be cut for the mass of region of $p\pi^{+}$ to be around $\Delta^{++}(1232)$ whose next excited state $\Delta(1600)$ is very far from it. So we will discuss $\Delta^{++}\pi^{-}$ instead of $p\pi^{+}\pi^{-}$ for the convenience. There are fruitful experimental data of the scatterings of $N\pi\to N\pi$ and $\Delta\pi$. Scientists have provided the real and imaginary parts of the partial-wave amplitudes of these scatterings, as an opne source in the website, such as the SAID program SAID . The details of the partial-wave amplitudes can be found in the website. This is most important point of this work. The long-distance contributions are usually very difficult to be calculated due to the non-perturbative property. It would significantly reduce the theoretical uncertainty if directly use the experimental data of scatterings. Besides, there are a lot of excited $N^{*}$ states whose masses and widths are of large uncertainties. The using of data of $N\pi$ scatterings would avoid the complex and unclear information of the resonant states. The real and imaginary parts of the partial-wave amplitudes contribute the strong phases which are necessary for CPV.
Besides, the isospin of the $N\pi$ system can only be 1/2, since the initial state $\Lambda_{b}$ is isospin singlet, and the weak transition of $b\to u$ changes the isospin by $\Delta I=1/2$. Therefore, only the partial-wave amplitudes with isospin 1/2 contribute to the $N\pi$ scatterings in our case. We will use the same symbols as the SAID program to represent the partial-wave amplitudes, $L_{2I,2J}$. $L$ is the angular momentum between $N$ and $\pi$, while $I$ and $J$ are the isospin and total spin of the system. For example, $D_{13}$ represents a partial-wave amplitude with $L=2$, $I=1/2$ and $J=3/2$. In fact, the spin-parity quantum states and the partial-wave amplitudes are one-to-one corresponding to each other, such as $1/2^{+}:P_{11}$, $1/2^{-}:S_{11}$, $3/2^{+}:P_{13}$, $3/2^{-}:D_{13}$, etc.
Figure 2: Angular definition of $\Lambda_{b}^{0}\to(p\pi^{0})h^{-}$ and $(\Delta^{++}\pi^{-})h^{-}$ with $h=\pi$ or $K$.
Now we can express the full decay amplitude of $\Lambda_{b}\to(\mathcal{B}M)h^{-}$ with $\mathcal{B}=p(\Delta^{++})$ and $M=\pi^{0}(\pi^{-})$ as
$\mathcal{A}(\Lambda_{b}\to(\mathcal{B}M)h^{-})$
$=$ $\lambda_{u}f_{h}\bar{u}_{N\pi}\bigg{[}a_{1}\left(P_{11}f_{1}^{1/2^{+}}-S_{11}f_{1}^{1/2^{-}}+\cdots\right)m_{-}$
$~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+a_{1}\left(P_{11}g_{1}^{1/2^{+}}-S_{11}g_{1}^{1/2^{-}}+\cdots\right)m_{+}\gamma_{5}\bigg{]}u_{\Lambda_{b}}$
$+$ $\lambda_{t}f_{h}\bar{u}_{N\pi}\bigg{[}\left(a_{46+}P_{11}f_{1}^{1/2^{+}}-a_{46-}S_{11}f_{1}^{1/2^{-}}+\cdots\right)m_{-}$
$~{}~{}~{}~{}~{}~{}~{}~{}+\left(a_{46-}P_{11}g_{1}^{1/2^{+}}-a_{46+}S_{11}g_{1}^{1/2^{-}}+\cdots\right)m_{+}\gamma_{5}\bigg{]}u_{\Lambda_{b}}$ (2)
where $\lambda_{i}=V_{ib}V_{iq}^{*}$ with $i=u,t$ and $q=d,s$, $f_{h}$ is the decay constant of $h=\pi$ or $K$, $a_{46\pm}=a_{4}\pm R_{h}a_{6}$, $R_{h}=2m_{h}^{2}/(m_{b}(m_{u}+m_{q}))$, and $a_{1,4,6}$ are the effective Wilson coefficients in the weak Hamiltonian Buchalla:1995vs ; Ali:1998eb , $m_{\pm}=m_{\Lambda_{b}}\pm m_{N\pi}$.
It can be seen from the above formula that the interference effects in the tree and penguin contributions are different. The above difference comes from the signs of the $a_{6}$ terms, compared to those of the $a_{1}$ and $a_{4}$ terms. In the four-fermion operators of $O_{1-4}$, quarks are all left-handed due to the weak interaction $(\bar{u}b)_{V-A}=\bar{u}_{L}\gamma^{\mu}b_{L}$. On the contrary, the $u$ quark from the $b$ transitions is right-handed with the operators of $O_{5,6}$, $(\bar{u}b)_{S-P}=\bar{u}_{R}b_{L}$. The different chiralities of $u$ quark result in different helicities of the $N\pi$ system, and hence different interference effects of partial-wave amplitudes. In the end, they contribute to the difference of strong phases for CPV.
Benefited by the partial-wave amplitudes, it can be easily studied for the angular analysis on $\Lambda_{b}$ decays. The angular distributions are illustrated in Fig. 2 for $\Lambda_{b}^{0}\to(p\pi^{0})h^{-}$ and $(\Delta^{++}\pi^{-})h^{-}$. The details of angular distributions can be easily obtained by the helicity amplitudes. The angle $\theta$ is defined between the directions of momentum of the final baryon $p(\Delta^{++})$ in the rest frame of $N\pi$ system and the momentum of $N\pi$ in the rest frame of $\Lambda_{b}^{0}$.
Numerical Discussions.
Table 1: Global CPV of $\Lambda_{b}$ decays stemming from $N\pi$ scatterings.
[…]
$N\pi\to\Delta^{++}\pi^{-}$ $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-}$ 6.0
$m_{N\pi}\in[1.2,1.9]$GeV $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})\pi^{-}$-4.4
$N\pi\to p\pi^{0}$ $\Lambda_{b}^{0}\to(p\pi^{0})K^{-}$ 5.8
$m_{N\pi}\in[1.1,2.5]$GeV $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$-3.9
Based on the above CPV mechanism and the data of $N\pi$ scatterings, we can predict the global and regional CP asymmetries of $\Lambda_{b}^{0}\to(p\pi^{0})h^{-}$ and $(\Delta^{++}\pi^{-})h^{-}$. The inputs of effective Wilson coefficients are given in Ali:1998eb . The inputs of $N\pi$ scattering amplitudes are model-independent and given in SAID . Since the data of scatterings of $N\pi\to N\pi$ and $\Delta\pi$ are very precise, we neglect the uncertainties of scattering amplitudes. Little is known about the form factors of $\Lambda_{b}\to N\pi$ at the current stage. But it is similar to the transition form factors of $B\to\pi\pi$ which have been studied by the heavy meson chiral perturbation theory Lee:1992ih ; Fajfer:1998yc ; Cheng:2007si ; Cheng:2020ipp and the light-cone QCD sum rules Cheng:2017smj ; Cheng:2017sfk . Therefore the form factors of $\Lambda_{b}\to N\pi$ can be calculated using the proven methods in principle. At the current stage, we will take some simple and typical values of these form factors, in order to reveal the new CPV mechanism. The point here is to show that large regional CPV can be generated in this mechanism, but not do a very precise prediction. Besides, the absolute values of form factors are cancelled in the ratio of CPV definition. So CPV relies only on the relative values of form factors. In the heavy quark limit, $f_{1}^{i}(0)=g_{1}^{i}(0)$. It is known that the heavy quark symmetry violating effect in the heavy-light baryonic form factors could be about $20\%$ Han:2022srw ; Huang:2022lfr . Therefore, we will consider a heavy-quark symmetry violating effect as $20\%$ without loss of generality. Besides, the mass regions of $p\pi^{0}$ and $\Delta^{++}\pi^{-}$ are taken as what SAID program provides, i.e. $m_{p\pi^{0}}\in[1.1,2.5]$GeV and $m_{\Delta^{++}\pi^{-}}\in[1.2,1.9]$GeV.
We can get the numerical results using the above inputs. Firstly, the global CPV in the mass regions of the $N\pi$ scattering system are given in Table. 1. It can be found that the magnitudes of global CPV of these processes are around $(4-6)\%$. Such values of CPV could be measured by LHCb for $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ and $(p\pi^{+}\pi^{-})\pi^{-}$ whose data samples are large currently with all the charged final-state particles.
Figure 3: CPV distributions of the Dalitz plots. The four plots are $\Lambda^{0}_{b}\to(p\pi^{0})\pi^{-}$, $\Lambda^{0}_{b}\to(p\pi^{0})K^{-}$, $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})\pi^{-}$ and $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})K^{-}$, respectively.
The regional CPV in the Dalitz plots are even more interesting. Every partial-wave amplitude corresponds to different angular distribution. Then the interference effects in the different regions in the Dalitz plots would induce different CP asymmetries. The numerical results of regional CPV are given in Fig. 3. It could found that the CPV could be as large as $20\%$ or even larger in some regions. This is large enough for the observation in experiments.
One more interesting property in the Dalitz plot is that there exists large regions with CPV of positive or negative values. This would be helpful for the experimental measurements. For example, it can be divided into two bins with $\cos\theta>0$ and $\cos\theta<0$, or say roughly $m_{hh}>1.5$GeV and $m_{hh}<1.5$GeV with $h=\pi$ or $K$, in all the four processes. The CPV in each bin could be large enough for experimental measurements, with large enough data samples.
From the Dalitz plots, it could be clearly seen that the regional CPV depend on $\cos\theta$ significantly. This is reasonable that the interference betwee different partial-wave amplitudes induce the dependence of CPV on $\cos\theta$. In order to see the dependence, it is very convenient to express the differential decay width on $\cos\theta$. The Fourier distributions after integrating out scattering energy are
$\frac{d\Gamma}{d\cos\theta}\propto\sum^{n=5}_{n=0}\mathcal{C}_{n}\cos n\theta$ (3)
with $\mathcal{C}_{n}=(1,-0.12,0.17,-0.02,0.003,0.02)$ for $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})K^{-}$, $(1,-0.29,0.14,-0.04,-0.05,0.02)$ for $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})\pi^{-}$, (1,1.3,0.7,0.8,0.03,0.13,(1,-1.3,0.7,-0.8,-0.03,-0.13,( 1 , - 1.3 , 0.7 , - 0.8 , - 0.03 , - 0.13 , 0.004)-0.004)- 0.004 ) for $\Lambda_{b}^{0}\to(p\pi^{0})K^{-}$, and (1,4.3,2.6,1.3,(1,-4.3,2.6,-1.3,( 1 , - 4.3 , 2.6 , - 1.3 , 0.4,0.2,0.02)0.4,-0.2,0.02)0.4 , - 0.2 , 0.02 ) for $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$, respectively. It can be found that the first two moments of $\cos\theta$ are large, so that the CPV in the Dalitz plots are basically caused by $\cos\theta$.
Figure 4: CPV of Fourier moments of $\cos(n\theta)$. The four plots are $\Lambda^{0}_{b}\to(p\pi^{0})\pi^{-}$, $\Lambda^{0}_{b}\to(p\pi^{0})K^{-}$, $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})\pi^{-}$ and $\Lambda^{0}_{b}\to(\Delta^{++}\pi^{-})K^{-}$, respectively.
So then the large regional CPV could also be measured by the moments of $\cos(n\theta)$. It can be easily realized by the experiments using the differential width convoluted with the moments of $\cos(n\theta)$. The results of CPV of Fourier moments are given in Fig. 4. The interesting result is that the CPV of the first moment $\cos\theta$ could be as large as $40\%$ in $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-}$. Considering the large component of the $\cos\theta$ term in the Eq. (3) for $\Lambda_{b}^{0}\to(\Delta^{++}\pi^{-})K^{-}$, it has a large possibility to observe the CPV of the $\cos\theta$ moment.
In the end, we strongly suggest to measure the CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ in the mass region of $m_{p\pi\pi}\in[1.2,1,9]$GeV, considering that $\Lambda_{b}^{0}\to p\pi^{+}\pi^{-}K^{-}$ has the largest data sample of charmless $b$-baryon decays. It could be in three ways. Firstly, the global CPV with the value of $6\%$ could be reached with all the data sample in the considered mass region. Secondly, the CPV in the Dalitz plot can be measured by two bins of $\cos\theta>0$ and $\cos\theta<0$, with the regional CPV of $(10-20)\%$. Thirdly, it is also realizable to measure the CPV of moment of $\cos\theta$ with a value as large as $40\%$.
Conclusions.— We propose a new CPV mechanism of baryon decays which stems from $N\pi$ scatterings. The weak phases come from the tree and penguin operators, while the strong phases are from the $N\pi$ scatterings. The most important advantage is that the fruitful data of $N\pi$ scatterings with precise partial-wave amplitudes could help us do model-independent analysis on the strong phases of $\Lambda_{b}$ decays. This is beneficial for our dynamical predictions on CPV of baryon decays. We study the CPV of $\Lambda_{b}^{0}\to(p\pi^{0})\pi^{-}$, $(p\pi^{0})K^{-}$, $(p\pi^{+}\pi^{-})\pi^{-}$ and $(p\pi^{+}\pi^{-})K^{-}$, via $N\pi\to p\pi^{0}$ and $N\pi\to\Delta^{++}\pi^{-}\to p\pi^{+}\pi^{-}$, respectively. It is found that the global CPV of these processes are at the order of several percents. The regional CPV of Dalitz plots could reach as large as $20\%$. The CPV of the moment $\cos\theta$ could be even larger. Such remarkable CPV have a large possibility to be observed in experiments. We strongly suggest to measure the CPV of $\Lambda_{b}^{0}\to(p\pi^{+}\pi^{-})K^{-}$ via either the global CPV or regional CPV or CPV of moment of $\cos\theta$ to achieve the first observation of baryon CPV.
Acknowledgement.—The authors are very grateful to Yanxi Zhang for the fruitful discussions on the experimental measurements at LHCb, and also to Xiaorui Lv, Wenbin Qian and Yangheng Zheng for the discussions on the prospects of experimental measurements. The authors are also grateful to Dan Guo and Igor Strakovsky for the introductions on the data of $N\pi$ scatterings from the SAID program SAID , and to Hai-Yang Cheng for the discussions on three-body $B$-meson decays. This work is supported in part by Natural Science Foundation of China under grant No. 12335003, and by the Fundamental Research Funds for the Central Universities under No. lzujbky-2024-oy02.